OpenStudy (crazynerd32):

A line has a slope of 4 and passes through the point (1, -1).

OpenStudy (oswaldmurphy):

Whats the question :/

OpenStudy (crazynerd32):

That's all it gave me so I think I have to create a equation.

OpenStudy (oswaldmurphy):

y = mx + b. M = slope.

OpenStudy (crazynerd32):

I know that but it has to pass through (1,-1,).

OpenStudy (oswaldmurphy):

www.desmos.com/calculator

OpenStudy (oswaldmurphy):

Thats a way you can check to look to see if it makes a line.

OpenStudy (oswaldmurphy):

ok since there is the point, which one is the x? (that will represent x in y = mx + b)

OpenStudy (crazynerd32):

1

OpenStudy (crazynerd32):

The slope is 4x

OpenStudy (oswaldmurphy):

y = 4(1) + b

OpenStudy (oswaldmurphy):

NEVER MIND IM AN IDIOT

OpenStudy (oswaldmurphy):

y - y1 = m(x - x1). replace this, do not solve.

OpenStudy (crazynerd32):

-1-1(-1)=4(4x-4(4) << is that how I will put it in.

OpenStudy (oswaldmurphy):

No, you do not have y or x. You have y1 and x1. You will keep y and x the same, replace the rest :D

OpenStudy (crazynerd32):

O ok.

OpenStudy (oswaldmurphy):

I will just give you step by step instruction. y - y1 first. y1 is -1 so y + 1 (the double negative cancles and becomes a positive. So you have y + 1 = m(x-x1) m = 4 as you already found out so y + 1 = 4(x - x1). Now x1 = 1 so your answer is: y + 1 = 4(x - 1)

OpenStudy (crazynerd32):

I get it now thx I will medal you and fan if you like.

OpenStudy (oswaldmurphy):

Sure, thank you. I just learned this too so I like practicing and teaching at the same time :D